OpenStudy (anonymous):

an objects motion is described by the equation d=4sin(pi t). the displacement d, is measured in meters. the time t is measure in seconds. 1. what is the objects position at t=0? be sure to include appropriate units 2. what is the objects maximum displacement from its resting position? be sure to include appropriate unit. 3. how much time is required for one oscillation. 4. what is the frequency of this motion? 5. what will the height of the object be at t=1.75 seconds?

OpenStudy (anonymous):

Hey sexy.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

1) at t = 0 d = 0 2) maximum displacement is nothing but maximum value of 4 sin(pie*t) = 4 3) One oscillation means total angle rotates by 2*pie ... at t=0 angle =0, at t = 2 angle = 2*pie .. hence t = 2 4) frequency = 1/t = 1/2 = 0.5 5) height = this depends on what path is the object moving .. is this a circular path or a straight line path .. and displacement is wrt a point or a line ?